OpenStudy (math22):

Tisha wants to go to Courpus Christi this weekend. She plans drive 50 miles per hour . Her distance can be calculated by using the equation d=50t , where d is the distance she travels and t is the time in hours . If Corpous Christi is 125 miles away , how long will it take her to get there?

OpenStudy (alfie):

50miles: 1hour = 125miles : x hours

OpenStudy (alfie):

Which is 125/50 = 2,5hours. You can do it by your formula with: d = 50t 125=50t t= 125/50 = 2.5hours